# Can LLMs Predict Citation Intent? An Experimental Analysis of In-context Learning and Fine-tuning on Open LLMs

### v1

@@ Abstract
This work investigates the ability of open Large Language Models (LLMs) to predict citation intent through in-context learning and fine-tuning. Unlike traditional approaches that rely on pre-trained models like SciBERT, which require extensive domain-specific pretraining and specialized architectures, we demonstrate that general-purpose LLMs can be adapted to this task with minimal task-specific data. We evaluate twelve model variations across five prominent open LLM families using zero, one, few, and many-shot prompting to assess performance across scenarios. Our experimental study identifies the top-performing model through extensive experimentation of in-context learning-related parameters, which we fine-tune to further enhance task performance. The results highlight the strengths and limitations of LLMs in recognizing citation intents, providing valuable insights for model selection and prompt engineering. Additionally, we make our end-to-end evaluation framework and models openly available for future use.
Can LLMs Predict Citation Intent? An Experimental Analysis of In-context Learning and Fine-tuning on Open LLMs
Paris Koloveas IMSI, ATHENA RC / Athens, GR University of the Peloponnese / Tripolis, GR pkoloveas@athenarc.gr                        Serafeim Chatzopoulos IMSI, ATHENA RC / Athens, GR schatz@athenarc.gr
Thanasis Vergoulis IMSI, ATHENA RC / Athens, GR vergoulis@athenarc.gr                        Christos Tryfonopoulos University of the Peloponnese / Tripolis, GR trifon@uop.gr
@@ 1 Introduction
Citations are references of research articles to external sources of information included to support claims, provide context, criticize, or acknowledge prior work. Although their primary function is to inform and redirect the reader, citations are also frequently used for other purposes, such as serving as proxies of scientific impact in various types of analysis (Kanellos et al., 2021). In such use cases, understanding the exact intent of a citation is crucial, as not all types of citations should be considered. For instance, while measuring the impact of an article, citations that criticize the work should generally not be considered as contributing to its impact.
Predicting citation intent based on its context (i.e., the sentences in the manuscript that accompany the citation) and other related information, has become an important classification problem to support the aforementioned use cases. Existing approaches have traditionally relied on linguistic features (Jurgens et al., 2018), machine learning methods (Teufel et al., 2006), and, in recent years, on pre-trained language models (PLMs) such as SciBERT (Beltagy et al., 2019) that require large scientific datasets such as (Lo et al., 2020) and task-specific architectures. In contrast, this work is the first to explore the potential of open, general-purpose Large Language Models (LLMs) to accurately identify citation intent without using any scientific-trained PLMs, evaluating their effectiveness through in-context learning and fine-tuning on minimal task-specific data.
Large Language Models (LLMs) are advanced natural language processing systems trained on extensive text corpora to perform a wide range of language tasks. Unlike traditional pre-trained language models (PLMs), LLMs are general-purpose models capable of adapting to new tasks with minimal additional training. Their ability to process and generate coherent text across diverse contexts makes them particularly suitable for tasks like citation intent classification, where understanding nuanced language patterns is essential.
In-context learning is a paradigm that allows language models to learn tasks given only a few examples in the form of demonstration (Brown et al., 2020; Dong et al., 2024). In-context learning is particularly suited to citation intent classification, as it allows the models to leverage their contextual understanding of language to make accurate predictions.
@@ 2.1 Models
For our experiments, we selected the Instruction-tuned versions of five prominent open-weight model families: LLaMA (Grattafiori et al., 2024), Mistral (Jiang et al., 2023), Phi (Abdin et al., 2024), Gemma (Gemma et al., 2024), and Qwen (Bai et al., 2023; Qwen et al., 2025). A key challenge was performing the experiments on commodity hardware, both due to computational limitations and to demonstrate that citation intent classification can be executed efficiently with limited resources. This influenced the selection of models, which range in size from small ($1$B parameters) to medium-sized ($32$B parameters).
@@
To reduce the memory footprint and computational requirements of our evaluation, we utilize the 8-bit (Q8) quantized versions of the models. This approach significantly reduces memory usage without compromising performance. Quantization involves converting model parameters from 16-bit floating-point precision to 8-bit integers, enabling more efficient computation while preserving expressive power and accuracy (Jurafsky and Martin, 2025).
The model variations used for our evaluation were the following (a more detailed view is available in Appendix C):
@@ 2.2 Datasets
  • SciCite (Cohan et al., 2019) consists of approximately $11,000$ citation strings annotated with three classes: Background Information, Method, and Results Comparison.
  • ACL-ARC (Jurgens et al., 2018) contains 2,000 citation strings classified into six categories: Background, Motivation, Uses, Extends, Compares or Contrasts, and Future.
These two datasets are widely used in citation intent classification research. The SciCite dataset is larger and more diverse, while the ACL-ARC dataset offers a more granular classification scheme.
@@ 2.3 Configuration Parameters
In this section, we describe the configuration parameters of the examined models, including prompting methods, system prompts, query templates, example methods, and temperature settings.
2.3.1 In-context Learning Methods
@@
To evaluate model performance, we applied four prompting methods (PM) by following the In-Context Learning paradigm (Brown et al., 2020): Zero-shot (no examples), One-shot (a single example per class), Few-shot ($5$ examples per class), and Many-shot ($10$ examples per class).
@@
The aforementioned prompting methods were selected to evaluate whether model performance improves as the number of examples increases, and to identify any saturation point where additional examples yield diminishing or negative returns. This aligns with prior literature on in-context learning methods, where these specific configurations have been extensively studied (Dong et al., 2024). Incorporating examples directly into the prompts allows models to better understand the task and the expected output format, which is especially important for citation intent classification, where the citation context is key to determining the correct label.
2.3.2 System Prompts
System prompts (SP) play a critical role in our experimental design, as they set the context and expectations for the model’s behavior. They guide the model’s attention on citation intent classification, providing task-specific context, class definitions, and guidelines to ensure outputs are aligned with the evaluation requirements.
We explored three distinct variations of system prompts. The first (SP1) was a simple, intuitive instruction, serving as a natural starting point for introducing the task. The second prompt (SP2) adopted a structured approach inspired by the CO-STAR framework (Teo, 2023), which organizes prompts into six key components. In particular, Context provides background information to help the model understand the scenario, while Objective clearly defines the task to direct its focus. Style, Tone, and Audience shape the model’s writing style, sentiment, and intended readership, respectively. Finally, Response defines the expected format to ensure clarity and downstream processing.
Since our task requires a single citation label as output, we excluded stylistic components (i.e., Style, Tone, and Audience), and incorporated detailed class definitions in the Response section. The third variation (SP3) refined the structured prompt by explicitly reiterating the expected labels in the Response section; this aims to further clarify the task, and enhance the consistency of the model’s outputs. All the prompts can be found in Appendix A.
2.3.3 Query Templates
Query templates (QT) define the structure in which citation sentences are presented to the model, both as examples during prompting and as queries during evaluation. Initially, we adopted a Simple Query template: the citation sentence was followed by “Class:”, either pre-filled with the correct class (in examples) or left blank (in queries). However, we observed that some models struggled to align with the expectation of responding solely with the class labels, leading to inconsistent performance and increased variance. To address this, we introduced a Multiple-choice Query template, where, following the citation sentence, the model is explicitly asked to identify the most likely citation intent, with the possible class labels presented as multiple-choice options. While this approach increases the token count per query, it demonstrates significant performance gains, as discussed in Section 3.2.
Dataset Rank Model Overall Zero-Shot One-Shot Few-Shot Many-Shot
SciCite 1 Qwen 2.5 – 14B 112 20 37 27 28
2 Mistral Nemo – 12B 22 0 0 12 10
3 Gemma 2 – 9B 4 0 3 1 0
4 Gemma 2 – 27B 2 0 0 0 2
ACL-ARC 1 Qwen 2.5 – 14B 24 4 4 8 8
2 Gemma 2 – 27B 4 0 4 0 0
Table 1: Model ranking based on Best-Performing Count (Overall & By Prompting Method).
Dataset Rank Model PM SP QT EM T F1-Score
SciCite 1 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 0.5 78.27
2 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 1.0 78.39
3 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 0.2 78.34
4 Qwen 2.5 – 14B Many-shot SP3 Multiple-Choice Roles 0.0 78.00
5 Qwen 2.5 – 14B Many-shot SP3 Multiple-Choice Roles 0.5 77.93
ACL-ARC 1 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 1.0 60.88
2 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 0.0 59.75
3 Qwen 2.5 – 14B Few-shot SP3 Multiple-Choice Roles 0.2 59.66
4 Qwen 2.5 – 14B One-shot SP3 Multiple-Choice Roles 0.5 64.87
5 Qwen 2.5 – 14B One-shot SP3 Multiple-Choice Roles 1.0 64.32
Table 2: Top-5 Configurations on SciCite and ACL-ARC (Q8).
2.3.4 Example Method Variations
@@
In our experiments, we explored two methods for presenting examples to the models (EM): Inline and Roles.
In the Inline approach, example citation sentences and their corresponding classes (as per the defined query templates) are provided directly within the system prompt. After this setup, the prompt then transitions to the evaluation phase, where the “user” provides a citation sentence without a corresponding class and the model, acting as the “assistant”, predicts the correct class.
The Roles method, on the other hand, simulates a conversational exchange between the “user” and “assistant”. Each example follows a back-and-forth interaction, with the “user” providing a citation sentence and the “assistant” responding with the correct class. After several such examples, the interaction moves to the evaluation phase, where the “user” presents a new citation sentence, and the model predicts its class.
2.3.5 Temperature
@@
Temperature (T) is a hyperparameter that controls the randomness or creativity of a language model’s outputs by adjusting the probability distribution of possible next tokens (Jurafsky and Martin, 2025). Lower temperatures (i.e., close to 00) correspond to greedy decoding, where the model deterministically selects the most probable token at each step, while higher temperatures (closer to $1$) introduce greater variability by allowing the model to sample from a broader range of options.
@@
For classification tasks, we aim for the model to output the most probable class label. Therefore, we use a temperature of 00 as the baseline, ensuring fully deterministic predictions. To explore how controlled randomness affects the model’s behavior and classification performance, we also evaluate higher temperatures, such as 0.2, 0.5, and 1.0. A temperature of 0.2 introduces a small degree of randomness but still heavily favors the most probable answer, while 0.5 strikes a balance between randomness and determinism. At 1.0, randomness is maximized, allowing us to assess whether excessive stochasticity degrades performance.
@@ 3 Experimental Results and Analysis
This section presents the results of our experiments, focusing on the performance of models across various configurations and parameter settings. We analyze the best-performing models, evaluate their rankings, and explore how specific parameters influence performance.
@@ 3.1 Model Performance over Configurations
3.1.1 Best-Performer Evaluation
To identify the most effective models across configurations, we initially conducted a Best-Performer Evaluation. In particular, for each configuration, we examined a metrics table containing precision, recall, F1-score, and accuracy for all models. The model with the highest F1-score in each configuration was selected as the best-performing model – in case of a tie, we used Accuracy as the deciding factor.
The aggregated results of our evaluation are presented in Table 1; additionally, Table 2 lists the top-$5$ best-performing configurations on each dataset. It is evident that Qwen 2.5 14B was the most dominant model on both datasets, significantly outperforming others across all prompting methods.
Rank Model Score
1 Qwen 2.5 – 14B 124.164
2 Mistral Nemo – 12B 59.442
3 Gemma 2 – 27B 48.421
4 Gemma 2 – 9B 47.904
5 Qwen 2 – 7B 31.927
6 LLaMA 3.1 – 8B 23.591
7 LLaMA 3 – 8B 21.535
8 Phi 3.5 Mini – 3.8B 18.869
9 Phi 3 Medium – 14B 16.877
10 Gemma 2 - 2B 15.467
11 LLaMA 3.2 - 3B 14.603
12 LLaMA 3.2 - 1B 11.620
Table 3: Model ranking based on RRF on SciCite.
3.1.2 Ranked Evaluation
While the Best-Performer Evaluation provided insights into the most dominant models across configurations, it lacked the ability to capture nuances among high-performing models. For instance, a model that consistently ranked second in multiple configurations would not be reflected in that approach. To address this limitation, we adopted a ranking methodology inspired by Reciprocal Rank Fusion (RRF) (Cormack et al., 2009). This approach allowed us to evaluate all $12$ models considering their relative performance across different configurations.

### v2

@@ Abstract
This work investigates the ability of open Large Language Models (LLMs) to predict citation intent through in-context learning and fine-tuning. Unlike traditional approaches relying on domain-specific pre-trained models like SciBERT, we demonstrate that general-purpose LLMs can be adapted to this task with minimal task-specific data. We evaluate twelve model variations across five prominent open LLM families using zero-, one-, few-, and many-shot prompting. Our experimental study identifies the top-performing model and prompting parameters through extensive in-context learning experiments. We then demonstrate the significant impact of task-specific adaptation by fine-tuning this model, achieving a relative F1-score improvement of 8% on the SciCite dataset and 4.3% on the ACL-ARC dataset compared to the instruction-tuned baseline. These findings provide valuable insights for model selection and prompt engineering. Additionally, we make our end-to-end evaluation framework and models openly available for future use.
@@ 1 Introduction
Citations are references of research articles to external sources of information included to support claims, provide context, criticize, or acknowledge prior work. Although their primary function is to inform and redirect the reader, citations are also frequently used for other purposes, such as serving as proxies of scientific impact in various types of analysis [18]. In such cases, understanding the exact intent of a citation is crucial, as not all types of citations should be considered. For instance, while measuring the impact of an article, citations that criticize the work should generally not be considered as contributing to its impact.
Predicting citation intent based on its context (i.e., the sentences in the manuscript that accompany the citation) and other related information, has become an important classification problem to support the aforementioned use cases. Existing approaches have traditionally relied on linguistic features [17], machine learning methods [32], and, in recent years, on domain-specific pre-trained language models (PLMs), such as SciBERT [3], that require large scientific datasets such as [21] and task-specific architectures [^1] In this paper, when referring to traditional PLMs in this context, we primarily mean models based on the encoder-only Transformer architecture, such as BERT and its derivatives (e.g., SciBERT), which were commonly fine-tuned for specific classification tasks. This distinguishes them from the general-purpose Large Language Models (LLMs) evaluated in this work, which typically utilize decoder-only or encoder-decoder Transformer architectures and are often leveraged for their generative and in-context learning capabilities.. In contrast, this work is the first to explore the potential of open, general-purpose Large Language Models (LLMs) – which, despite potentially encountering scientific text during their broad pre-training, are not specifically optimized for the scientific domain like SciBERT – to accurately identify citation intent, evaluating their effectiveness through in-context learning and fine-tuning on minimal task-specific data.
Large Language Models are advanced natural language processing systems trained on extensive text corpora to perform a wide range of language tasks. Unlike traditional pre-trained language models, LLMs are general-purpose models capable of adapting to new tasks with minimal additional training. Their ability to process and generate coherent text across diverse contexts makes them particularly suitable for tasks like citation intent classification, where understanding nuanced language patterns is essential.
In-context learning is a paradigm that allows LLMs to learn tasks given only a few examples in the form of demonstration [4, 8]. In-context learning is particularly suited to citation intent classification, as it allows the models to leverage their contextual understanding of language to make accurate predictions.
@@ 2.1 Models
For our experiments, we selected the Instruction-tuned versions of five prominent open-weight model families: LLaMA [11], Mistral [15], Phi [1], Gemma [10], and Qwen [2, 26]. A key challenge was performing the experiments on commodity hardware, both due to computational limitations and to demonstrate that citation intent classification can be executed efficiently with limited resources. This influenced the selection of models, which range in size from small ($1$B parameters) to medium-sized ($27$B parameters).
@@
To reduce the memory footprint and computational requirements of our evaluation, we utilize the 8-bit (Q8) quantized versions of the models. This approach significantly reduces memory usage without compromising performance. Quantization involves converting model parameters from higher numerical precisions (e.g. 16-bit floating-point) to formats of lower numerical precision (e.g. 8-bit integers). This process enables more efficient computation and reduced memory usage while largely preserving model performance and expressive power [16].
The model variations used for our evaluation were the following:
@@ 2.2 Datasets
  • SciCite [5] consists of $11,021$ citation strings annotated with three classes: Background Information, Method, and Results Comparison.
  • ACL-ARC [17] contains $1,941$ citation strings split into six categories:
    Background, Motivation, Uses, Extends, Compares or Contrasts, and
    Future.
These two datasets are widely used in citation intent classification research. The SciCite dataset is larger and more diverse, with broad coverage across multiple scientific fields, namely, Computer Science, Medicine, Neuroscience, and Biochemistry, while the ACL-ARC dataset offers a more granular classification scheme, focused on Computational Linguistics. Accross our evaluation, we used the original splits of the datasets, which include training, validation, and test sets (75% - 8% - 17% for SciCite, and 87% - 7% - 6% for ACL-ARC).
@@ 2.3 Configuration Parameters
In this section, we describe the configuration parameters of the examined models, including in-context learning methods, system prompts, query templates, example methods, and temperature settings.
2.3.1 In-context Learning Methods (ILM).
@@
To evaluate model performance, we applied four prompting methods [^4] Throughout the paper, we use “In-context Learning Methods” and “Prompting Methods” interchangibly. by following the In-Context Learning paradigm [4]: Zero-shot (no examples), One-shot (a single example per class), Few-shot ($5$ examples per class), and Many-shot ($10$ examples per class).
@@
These prompting methods were selected to evaluate whether model performance improves as the number of examples increases, and to identify any saturation point where additional examples yield diminishing or negative returns. This aligns with prior literature on in-context learning methods, where these specific configurations have been extensively studied [8]. Incorporating examples directly into the prompts allows models to better understand the task and the expected output format, which is especially important for citation intent classification, where the citation context is key to determining the correct label.
Figure 1: Our System Prompts.
2.3.2 System Prompts (SP).
System prompts (SP) are foundational to our experimental design, establishing the context and expected behavior for the models in the citation intent classification task. They provide task-specific instructions, class definitions, and output guidelines. We developed and evaluated three distinct system prompt variations, shown in Figure 1.
  • SP1: This initial prompt offered a straightforward, intuitive instruction for the classification task.
  • SP2: This prompt introduced more structure, drawing inspiration from the CO-STAR framework [31]. We adapted relevant CO-STAR components by definint the task Context, the Objective of the model, and detailed Response Rules. Given that our task requires a single, specific class label output, the stylistic components of CO-STAR (Style, Tone, and Audience) were not incorporated.
  • SP3: Building upon SP2, this prompt further refined wording of the Objective and Response Rules. It also explicitly reiterated the expected labels in the Response Rules section, aiming to improve task clarity and consistency in the model’s outputs.
2.3.3 Query Templates (QT).
Query templates (QT) define the specific format used to present citation sentences to the model, whether as part of an in-context learning example or as the actual query requiring classification. We evaluated two template structures:
  • Simple Query: Initially, we used a basic format where the citation sentence was followed by “Class:”. For examples, the correct class was appended; for queries, this was left for the model to predict.
  • Multiple-choice Query: We observed that some models struggled to adhere to the strict class-label-only output expected with the Simple Query, resulting in inconsistent performance. To mitigate this, we introduced a more structured template. Here, after the citation sentence, the model is explicitly prompted to choose the most appropriate citation intent from a presented list of all possible class labels.
Although the Multiple-choice template increases the token count per instance, it yielded significant performance improvements (detailed in Section 3.2) by providing clearer guidance to the models.
2.3.4 Example Presentation Formats (EPF).
@@
In our experiments, we explored two formats for presenting in-context examples to the models: Inline and Conversational.
  • Inline: Example citation sentences and their corresponding classes (structured according to the defined query templates) are embedded directly and sequentially within the main prompt, immediately preceding the actual query sentence that the model needs to classify.
  • Conversational: Examples are structured to simulate a dialogue. Each example consists of a “user” turn providing a citation sentence and an “assistant” turn providing the correct class. This series of example turns is provided before the final “user” query, which presents the new citation sentence for the model to classify. This format is also commonly referred to as a conversational or turn-taking format.
2.3.5 Temperature (T).
@@
Temperature is a hyperparameter that controls the randomness or creativity of a language model’s outputs by adjusting the probability distribution of possible next tokens [16]. Lower temperatures (i.e., close to 00) correspond to greedy decoding, where the model deterministically selects the most probable token at each step, while higher temperatures (closer to $1$) introduce greater variability by allowing the model to sample from a broader range of options.
@@
For classification tasks, we aim for the model to output the most probable class label. Therefore, we use a temperature of 00 as the baseline, ensuring fully deterministic predictions. To explore how controlled randomness affects classification performance, we also evaluate higher temperatures, such as $0.2$, $0.5$, and $1.0$. A temperature of $0.2$ introduces a small degree of randomness but still heavily favors the most probable answer, while $0.5$ strikes a balance between randomness and determinism. At $1.0$, randomness is maximized, allowing us to assess whether excessive stochasticity degrades performance.
@@ 3 Experimental Results and Analysis
This section presents the results of our experiments, focusing on model performance across various configurations and parameter settings. An initial overview of the peak F1-score achieved by each of the twelve models on both the SciCite and ACL-ARC datasets is summarized in Table 1. The subsections that follow provide a detailed analysis of top-performing model configurations, evaluate overall model rankings, and explore the influence of specific parameters.
Table 1: Highest Performance by Model.
SciCite ACL-ARC
Rank Model F1-Score Rank Model F1-Score
1 Qwen 2.5 – 14B 78.33 1 Qwen 2.5 – 14B 63.68
2 Gemma 2 – 27B 77.86 2 Gemma 2 – 27B 58.95
3 Mistral Nemo – 12B 77.39 3 Gemma 2 – 9B 57.19
4 Gemma 2 – 9B 75.12 4 Qwen 2 – 7B 51.26
5 Phi 3 Medium – 14B 74.67 5 LLaMA 3.1 – 8B 48.45
6 LLaMA 3 – 8B 74.39 6 Mistral Nemo – 12B 48.11
7 Qwen 2 – 7B 72.89 7 Phi 3.5 Mini – 3.8B 43.74
8 LLaMA 3.1 – 8B 72.46 8 Phi 3 Medium - 14B 43.46
9 Gemma 2 - 2B 68.79 9 Gemma 2 - 2B 40.96
10 Phi 3.5 Mini – 3.8B 68.25 10 LLaMA 3.2 - 3B 40.07
11 LLaMA 3.2 - 3B 67.99 11 LLaMA 3 - 8B 38.06
12 LLaMA 3.2 - 1B 45.44 12 LLaMA 3.2 - 1B 24.60
@@ 3.1 Model Performance over Configurations
3.1.1 Best-Performer Evaluation.
To identify the most effective models across configurations, we initially conducted a Best-Performer Evaluation. In particular, for each configuration, we examined a metrics table containing precision, recall, F1-score, and accuracy for all models. The model with the highest F1-score [^7] All F1-scores reported in this paper are macro-averaged F1-scores (macro-F1). in each configuration was selected as the best-performing model – in case of a tie, we used Accuracy as the deciding factor.
The aggregated results of our evaluation are presented in Table 2. It is evident that Qwen 2.5 14B was the most dominant model on both datasets, significantly outperforming others across all prompting methods.
Table 2: Model Ranking based on Best-Performing Count.
Dataset Rank Model Overall Zero-Shot One-Shot Few-Shot Many-Shot
SciCite 1 Qwen 2.5 – 14B 125 24 42 28 31
2 Mistral Nemo – 12B 25 0 0 14 11
4 Gemma 2 – 27B 10 0 0 4 6
4 Gemma 2 – 9B 8 0 6 2 0
ACL-ARC 1 Qwen 2.5 – 14B 129 19 22 40 28
2 Gemma 2 – 27B 28 0 21 7 0
3 Gemma 2 – 9B 11 5 5 1 0
Table 3: RRF-based Model Ranking.
SciCite ACL-ARC
Rank Model Ranking Score Rank Model Ranking Score
1 Qwen 2.5 – 14B 144.163 1 Qwen 2.5 – 14B 146.496
2 Mistral Nemo – 12B 68.479 2 Gemma 2 – 27B 62.973
3 Gemma 2 – 27B 63.369 3 Gemma 2 – 9B 55.318
4 Gemma 2 – 9B 59.760 4 Qwen 2 – 7B 41.496
5 Qwen 2 – 7B 38.617 5 LLaMA 3.1 – 8B 40.844
6 LLaMA 3.1 – 8B 27.624 6 Phi 3.5 Mini – 3.8B 30.470
7 LLaMA 3 – 8B 26.159 7 Mistral Nemo – 12B 30.285
8 Phi 3.5 Mini – 3.8B 22.088 8 LLaMA 3.2 - 3B 23.984
9 Phi 3 Medium – 14B 19.370 9 Phi 3 Medium - 14B 21.269
10 Gemma 2 - 2B 19.253 10 LLaMA 3.2 - 1B 16.796
11 LLaMA 3.2 - 3B 18.286 11 Gemma 2 - 2B 16.534
12 LLaMA 3.2 - 1B 14.136 12 LLaMA 3 - 8B 16.359
3.1.2 Ranked Evaluation.
While the Best-Performer Evaluation provided insights into the most dominant models across configurations, it lacked the ability to capture nuances among high-performing models. For instance, a model that consistently ranked second in multiple configurations would not be reflected in that approach. To address this limitation, we adopted a ranking methodology inspired by Reciprocal Rank Fusion (RRF) [6]. This approach allowed us to evaluate all $12$ models considering their relative performance across different configurations.